OpenStudy (anonymous):

Help! Will Medal!! :) A system of equations is shown below: 5x - 5y = 10 3x - 2y = 2 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

I think I've figured out the some of the first part but I don't know where to go from there..

OpenStudy (anonymous):

You could subsitute\[3x-2y=2\]\[-2y -3x=2-3x\]\[-2y=2-3x\]\[-2y/-2=(2-3x)/-2\]\[y=(2-3x)/-2\]

OpenStudy (anonymous):

Do you know what to do next?

OpenStudy (anonymous):

No I'm confused. Could you write out the steps please I'm not sure how you got to each :(

OpenStudy (anonymous):

I did! I was just setting one of the equations equal to y

OpenStudy (anonymous):

I am not finished yet. I was seeing if you knew where to go from where I left off.

OpenStudy (anonymous):

I meant worded form. Like what did you sub in for that equation? And is your goal to isolate y?

OpenStudy (anonymous):

I isolated y first

OpenStudy (anonymous):

Do you remember how I got\[(2-3x)/-2\]For y?

OpenStudy (anonymous):

Well now since I know what y equals I can substitute y like this: \[ 5x - 5y = 10\] \[5x - 5((2-3x)/-2) = 10\]

OpenStudy (anonymous):

Are you following?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Alright, now we solve for x.

OpenStudy (anonymous):

Okay. Wait is this still only PART A?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

Okay. Would my answer be 6/5?

OpenStudy (anonymous):

For x

OpenStudy (anonymous):

Oh wait! I'm approaching A the wrong way.

OpenStudy (anonymous):

Sorry, it's like this because it is "Replace one equation with the sum of that equation and a multiple of the other."

OpenStudy (anonymous):

Nevermind

OpenStudy (anonymous):

:( Im confused

OpenStudy (anonymous):

Here is an archived question that might be able to help: http://openstudy.com/study#/updates/52eb025ce4b0d95ca6b33161

OpenStudy (anonymous):

This method is what I was doing before and I got: Solve the system by multiplying the top by 3 and the bottom by -5. So 15x-15y=30 -15x+10y=-10 This equals -5y=20 and so y=-4

OpenStudy (anonymous):

But I don't know if it's right or where I go from here :L

OpenStudy (anonymous):

Don't you have to multiply one by a negative though? Otherwise it doesn't cancel out?

OpenStudy (anonymous):

no need to as we are just isolating the y and disposing the x.

OpenStudy (anonymous):

give us a sec to alter it... jst a slight mistake

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

I'll write on on paper and scan it and post it.

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

sorry..the handwriting is bad

OpenStudy (anonymous):

It's lovely. And thank you. Part B? You might only have to explain it to me because I just don't understand what the question is asking of me.

OpenStudy (anonymous):

Just show that that your first pair of equations and your second pair of equations have the same solution by solving for y and x on both sets.

OpenStudy (anonymous):

I am not really sure as well. It could be that you have to substitute both "X" and "Y" into both 1a and 2a to prove that they have the same answer. So, you should get the same answers as the first pair of equations using the second pair of equations.

OpenStudy (anonymous):

Okay Thanks! I will fan you both

OpenStudy (anonymous):

anymore question?

OpenStudy (anonymous):

No but I'll let you know if I do! Thanks!